# Combining Evidence Across Filtrations

### v3

@@ 6 Characterizing adjusters as $\mathfrak{e}$-lifters
In the early works of Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011); Koolen and Vovk, (2014), an adjuster is defined as a function that turns the running maximum of a test supermartingale into a process that is upper-bounded by another test supermartingale. These works show that their definition is equivalent to the analytic definition of adjusters (11). We shall refer to their definition as game-theoretic because it views a test supermartingale as the wealth of a skeptic who bets against the null. In this view, an adjuster is a function that turns the running maximum of a skeptic’s wealth into a valid bet (i.e., the skeptic is not expected to make money if the data comes from the null). We include a game-theoretic interpretation of $\mathfrak{e}$-lifting in Section H.
@@ 6.1 Known characterization of adjusters for test supermartingales
We begin by reviewing the notion of “adjusters for test supermartingales.” This appears in each of the aforementioned early works, although the work of Dawid et al., 2011b is the most relevant to our setup. Their work (theorem 1) shows that an increasing function $\mathsf{A}:[1,\infty]\to[0,\infty]$ is an adjuster if and only if, for any probability distribution $P$ and for any test supermartingale $(M_{t})_{t\geq 0}$ for $P$ in some filtration $\mathbb{G}$, there exists another test supermartingale $(M_{t}^{\prime})_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the adjusted running maximum process $(\mathsf{A}(M_{t}^{*}))_{t\geq 0}$, that is,
@@ 6.2 The main characterization theorem
We now introduce the main result of this section, synthesizing the two characterizations of adjusters with Theorem 2. The takeaway is that all notions of adjusters coincide, including their characterization as $\mathfrak{e}$-lifters, as long as $\mathsf{A}$ is an increasing function applied to the running maximum of an e-process.
@@
[…]
@@
[…]
As discussed earlier, statements a and b are equivalent. Further, the $\mathfrak{e}$-lifting theorem says that a implies e. Notice also that e trivially implies d. It then suffices to show two nontrivial statements, which we prove in Section F.3, that connect the rest of the dots:
@@
Theorem 3 establishes an extent to which the use of an adjuster is necessary for lifting evidence: the implication e $\Rightarrow$ a says that, if $f$ is some increasing function that maps the running maximum of any e-process in $\mathbb{G}$ into an e-process in any $\mathbb{F}\supseteq\mathbb{G}$, for an arbitrary $\mathcal{P}$, then $f$ must be an adjuster. Further, even if we found an increasing function $f$ that maps, for any $\mathcal{P}$, the running maximum of any e-process in $\mathbb{G}$ into an e-process in some $\mathbb{F}\supseteq\mathbb{G}$, then we can still be assured that $f$ is an adjuster because $f$ satisfies condition d. Put together, we may conclude that, when it comes to mapping the running maximum of arbitrary e-processes to a valid e-process in some finer filtration, using an adjuster is necessary.
@@
We close with the note that the characterization of adjusters and $\mathfrak{e}$-lifters in Theorem 3 is specific to (increasing) functions that apply to the running maximum of e-processes. Thus, in theory, the theorem does not rule out the existence of an $\mathfrak{e}$-lifter that depends on, say, $\mathfrak{e}_{t}$ but not $\mathfrak{e}_{t}^{*}$, or other functions with entirely different forms. In Section D.2, we explore a class of generalized adjusters (Dawid et al., 2011a, ) that take both $\mathfrak{e}_{t}^{*}$ and $\mathfrak{e}_{t}$ as inputs, and we show how these are not $\mathfrak{e}$-lifters in general. Identifying alternate forms of $\mathfrak{e}$-lifters is left as future work.
@@ 7 Discussion
Theorem 2 implies that the answer is “no”: if we can find a powerful e-process in a sub-filtration, then we can always $\mathfrak{e}$-lift it into the data filtration using an adjuster. Since any admissible adjuster $\mathsf{A}$ satisfies $\lim_{\mathfrak{e}\to\infty}\mathsf{A}(\mathfrak{e})=\infty$ by definition (11), we can be assured that the e-process lifted by $\mathsf{A}$ is also powerful. In general, adjusting a test martingale in a sub-filtration yields an e-process but not a test martingale in the data filtration, so this result does not contradict prior results on how there exist no powerful test martingales for randomness or independence.
@@
Since adjusters sacrifice some amount of evidence, we may consider leveraging external randomization to recover the loss in statistical power. In Section I, we explore this idea in two ways; here, we briefly summarize our findings. The first approach is to borrow ideas from Ramdas and Manole, (2024), who utilize an independent $\mathsf{Unif}[0,1]$ random variable $U$ and use the threshold of $U/\alpha$ for testing at a stopping time. This “lift-then-randomize” approach is straightforwardly applicable to the adjusted e-process. The second approach is to build upon the ideas of Ignatiadis et al., (2024), who develop a randomized e-to-p calibration method for e-values. However, we find that the resulting “randomize-then-lift” approach does not generally yield anytime-valid procedures when applied to $\mathfrak{e}$-lifting.
@@ B.1 Problem setup
$\mathcal{H}_{0}^{\mathsf{IID}}:X_{1},X_{2},\dotsc\text{ are IID.}$. (28)
@@
$\mathcal{H}_{0}^{\mathsf{exch}}:X_{1},X_{2},\dotsc\text{ are exchangeable.}$. (29)
@@ B.3 Conformal test martingales and the “simple jumper” algorithm
=$=\begin{cases}U_{t}\cdot f_{1}(t)&\mbox{if }X_{t}=1;\\ f_{1}(t)+U_{t}\cdot f_{0}(t)&\mbox{if }X_{t}=0,\end{cases}$ (32)
where $f_{k}(t)=t^{-1}\cdot\#\{i\leq t:X_{i}=k\}$ denotes the empirical frequency of seeing $k\in\{0,1\}$ up to time $t$, and $U_{t}\sim\mathsf{Unif}[0,1]$ is an IID Uniform random variable that is independent of $\mathbb{F}$. Note that $s_{t}$ is based on summary statistics of the data up to time $t$ along with external randomization. Equation (31) is a general definition for conformal p-values, while equation (32) represents a simplification assuming that the nonconformity score is the identity function and that the data is binary. By Vovk et al., (2003); Vovk et al., 2022a , the conformal p-values $(s_{t})_{t\geq 1}$ are IID with $\mathsf{Unif}[0,1]$ under the null hypothesis $\mathcal{H}_{0}^{\mathsf{IID}}$. Despite its name, the conformal p-value is not a p-value in the traditional sense, and under changepoint alternatives, it deviates from $1/2$ to either 0 or 1.
@@
In practice, since a weighted average of e-processes in $\mathbb{G}$ is also an e-process in $\mathbb{G}$, we can take a mixture of $(\mathfrak{e}_{t}^{\mathsf{conf}}(\lambda))_{t\geq 0}$ defined using different values of $\lambda$. Vovk et al., (2021) defines the “simple jumper” conformal test martingale as
$\mathfrak{e}_{t}^{\mathsf{jump}}=w_{-1}\mathfrak{e}_{t}(-1)+w_{0}\mathfrak{e}_{t}(0)+w_{1}\mathfrak{e}_{t}(1),\quad\forall t\geq 0,$ (34)
for some fixed weights $w_{-1},w_{0},w_{1}\geq 0$ that sum to 1, or alternatively notice that (33) is also an e-process for any $\mathbb{G}$-predictable sequence $(\lambda_{t})_{t\geq 0}$ and use clever choices of the hyperparameter (e.g., Fedorova et al.,, 2012; Waudby-Smith and Ramdas,, 2024).
The e-process we use in Example 1 corresponds to the simple jumper with weights $w_{1}=1$ and $w_{0}=w_{-1}=0$. The mean of stopped e-values (right plot of Figure 1) also exceeds 1 when $w_{-1}=w_{1}=1/2$ and the stopping time is the first time when $k$ consecutive zeros or ones are observed, although the standard error over repeated runs is substantially larger. Vovk, 2021a ’s original version, with $w_{-1}=w_{0}=w_{1}=1/3$, does not empirically violate the stopping time validity in the data filtration $\mathbb{F}$, despite not having $\mathbb{F}$-anytime-validity in theory. For the simulations under the alternative hypothesis (Figures 5 and 9), and for the real data experiments (Section 4.2), we use the original version with $w_{-1}=w_{0}=w_{1}=1/3$.
We note that, in this paper, the specific choice of the simple jumper conformal test martingale is made solely for the clarity of exposition, and the naïve simple jumper algorithm is by no means the optimal version of the conformal test martingale. There exist other variants of conformal test martingales, including ones powerful against other types of alternatives; see Vovk et al., 2022a (, section 9). Yet, as referenced in the main text, even when combining different instantiations of conformal test martingales, the same issue of not having validity at data-dependent stopping times would still arise.
@@ B.4 Why we do not need to lift the UI e-process
Appendix C Testing a scale-invariant Gaussian mean
Here, we give another concrete example of an e-process defined in a coarser filtration $\mathbb{G}\subsetneq\mathbb{F}$ that is not valid at a $\mathbb{F}$-stopping time. This particular construction is due to Hendriksen et al., (2021); Pérez-Ortiz et al., (2024); we follow the exposition of the latter paper. The e-process described below can be combined with another e-process that is valid in $\mathbb{F}$, such as Wang and Ramdas, 2025a ’s e-process based on universal inference.
@@ Appendix C Testing a scale-invariant Gaussian mean
[…]
@@
For this problem, we can also consider combining different e-processes in an analogous manner to the case of testing randomness (Example 1). For instance, Wang and Ramdas, 2025a (, theorem 3.2) develop a different e-process for the same null hypothesis (assuming, for simplicity, that $\delta_{0}=0$). This e-process is based on UI (Wasserman et al.,, 2020), analogous to Ramdas et al., (2022)’s UI e-process for testing randomness (30), and it is valid under the data filtration $\mathbb{F}$. The e-process has the following closed-form expression:
@@
To combine the two e-processes $\mathfrak{e}^{\mathsf{MaxInv}}$ (37) and $\mathfrak{e}^{\mathsf{UI}}$ (39), we can employ the adjust-then-combine strategy (Corollary 2) to obtain an e-process in $\mathbb{F}$:
@@
Appendix D Other types of adjusters
D.1 “Zero” adjusters can be (even more) powerful $\mathfrak{e}$-lifters

### v4

@@ 6 Characterizing adjusters as $\mathfrak{e}$-lifters
In Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011); Koolen and Vovk, (2014), an adjuster is defined as a deterministic function that turns the running maximum of a test supermartingale into a process that is upper-bounded by another test supermartingale. These early works show that their definition is equivalent to the analytic definition of adjusters (11). We shall refer to their definition as game-theoretic because it views a test supermartingale as the wealth of a skeptic who bets against the null. In this view, an adjuster is a function that turns the running maximum of a skeptic’s wealth into a valid bet (i.e., the skeptic is not expected to make money if the data comes from the null). We include a game-theoretic interpretation of $\mathfrak{e}$-lifting in Section I.
@@ 6.1 Known characterization of adjusters for test supermartingales
We begin by reviewing the notion of “adjusters for test supermartingales.” This appears in each of the aforementioned early works, although the work of Dawid et al., 2011b is the most relevant here. Their work shows that an increasing function $\mathsf{A}:[1,\infty]\to[0,\infty]$ is an adjuster in the sense of definition (11) if and only if, for any probability distribution $P$ and for any test supermartingale $(M_{t})_{t\geq 0}$ for $P$ in some filtration $\mathbb{G}$, there exists another test supermartingale $(M_{t}^{\prime})_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the adjusted running maximum process $(\mathsf{A}(M_{t}^{*}))_{t\geq 0}$:
@@ 6.2 The main characterization theorem
We now introduce the main result of this section, synthesizing the two characterizations of adjusters with Theorem 2. The takeaway is that all notions of adjusters coincide, including their characterization as $\mathfrak{e}$-lifters, as long as $\mathsf{A}$ is an increasing, deterministic function applied to the running maximum of an e-process.
@@
    $\mathsf{A}$ is an adjuster in the sense of (11), that is, it satisfies $\int_{1}^{\infty}\frac{\mathsf{A}(\mathfrak{e})}{\mathfrak{e}^{2}}d\mathfrak{e}\leq 1$.
@@
Furthermore, $\mathsf{A}$ is admissible according to any one of these characterizations if and only if $\mathsf{A}$ is right-continuous, $\mathsf{A}(\infty)=\infty$, and $\int_{1}^{\infty}\frac{\mathsf{A}(\mathfrak{e})}{\mathfrak{e}^{2}}d\mathfrak{e}=1$.
As discussed earlier, statements a and b are equivalent. Further, the $\mathfrak{e}$-lifting theorem says that a implies e. Notice also that e trivially implies d. It then suffices to show two nontrivial statements, which we prove in Section G.3, that connect the rest of the dots:
@@
Theorem 3 establishes an extent to which the use of an adjuster is necessary for lifting evidence: the implication e $\Rightarrow$ a says that, if $f$ is some increasing function that maps the running maximum of any e-process in $\mathbb{G}$ into an e-process in any $\mathbb{F}\supseteq\mathbb{G}$, for an arbitrary $\mathcal{P}$, then $f$ must be an adjuster. Further, even if we found an increasing function $f$ such that, for any $\mathcal{P}$, its application to the running maximum of any e-process in $\mathbb{G}$ is an e-process in some $\mathbb{F}\supseteq\mathbb{G}$, then we can still be assured that $f$ is an adjuster because $f$ then satisfies condition d. Put together, we may conclude that, when it comes to mapping the running maximum of arbitrary e-processes to a valid e-process in some finer filtration, using an adjuster is necessary.
@@
We close with the note that the characterization of adjusters and $\mathfrak{e}$-lifters in Theorem 3 is specific to (increasing) functions that apply to the running maximum of e-processes. Thus, in theory, the theorem does not rule out the existence of an $\mathfrak{e}$-lifter that depends on, say, $\mathfrak{e}_{t}$ but not $\mathfrak{e}_{t}^{*}$, or other functions with entirely different forms. In Section E.2, we explore a class of generalized adjusters (Dawid et al., 2011a, ) that take both $\mathfrak{e}_{t}^{*}$ and $\mathfrak{e}_{t}$ as inputs, and we show how these are not $\mathfrak{e}$-lifters in general. There can also exist $\mathfrak{e}$-lifters for specific testing problems and for e-processes in particular filtrations, although their utility will be limited to such settings. Identifying alternate forms of $\mathfrak{e}$-lifters is left as future work.
@@ 7 Discussion
Theorem 2 implies that the answer is “no”: if we can find a powerful e-process in a sub-filtration, then we can always $\mathfrak{e}$-lift it into the data filtration using an adjuster. Since any admissible adjuster $\mathsf{A}$ satisfies $\lim_{\mathfrak{e}\to\infty}\mathsf{A}(\mathfrak{e})=\infty$ by definition (11), we can be assured that the e-process lifted by $\mathsf{A}$ is also powerful. We summarize this implication as a corollary.
Corollary 3.
Let $\mathcal{P}$ be a composite null and $\mathcal{Q}$ be a composite alternative. Suppose there exists a $\mathcal{Q}$-powerful e-process $(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$. Then, there always exists a $\mathcal{Q}$-powerful e-process for $\mathcal{P}$ in $\mathbb{F}$ (namely, $(\mathsf{A}(\mathfrak{e}_{t}^{*}))_{t\geq 0}$ for any admissible adjuster $\mathsf{A}$).
In general, adjusting a test martingale in a sub-filtration yields an e-process but not a test martingale in the data filtration, so this result does not contradict prior results on how there exist no powerful test martingales for randomness or independence.
@@
Is there a trade-off between the coarseness of the base filtration and the power of the lifted e-process?
At a first glance, it may appear that the coarseness of the base filtration $\mathbb{G}$, relative to the data filtration $\mathbb{F}$, must somehow affect the power of the adjusted e-process. However, because the adjuster acts as a one-time, uniform fee for achieving anytime-validity in the finest “relevant” filtration $\mathbb{F}$, adjustment itself is independent of how coarse $\mathbb{G}$ is. On the other hand, the coarseness of $\mathbb{G}$ can affect the power of the base e-process $(\mathfrak{e}_{t})_{t\geq 0}$ in different ways. As we already saw, coarsening the filtration by an “appropriate” amount is a generally effective strategy for designing powerful test martingales for highly composite nulls (e.g., exchangeability and independence). Yet, if $\mathbb{G}$ is too coarse, e.g., if it contains no information about the data, then any e-process in $\mathbb{G}$ would have to be powerless.
@@
Since adjusters sacrifice some amount of evidence, we may consider leveraging external randomization to recover the loss in statistical power. In Section J, we explore this idea in two ways; here, we briefly summarize our findings. The first approach is to borrow ideas from Ramdas and Manole, (2024), who utilize an independent $\mathsf{Unif}[0,1]$ random variable $U$ and use the threshold of $U/\alpha$ for testing at a stopping time. This “lift-then-randomize” approach is straightforwardly applicable to the adjusted e-process. The second approach is to build upon the ideas of Ignatiadis et al., (2024), who develop a randomized e-to-p calibration method for e-values. However, we find that the resulting “randomize-then-lift” approach does not generally yield anytime-valid procedures when applied to $\mathfrak{e}$-lifting.
@@ B.1 Problem setup
$\mathcal{H}_{0}^{\mathsf{IID}}:X_{1},X_{2},\dotsc\text{ are IID.}$ (28)
@@
$\mathcal{H}_{0}^{\mathsf{exch}}:X_{1},X_{2},\dotsc\text{ are exchangeable.}$ (29)
@@ B.3 Conformal test martingales and the “simple jumper” algorithm
$=\begin{cases}U_{t}\cdot f_{1}(t)&\mbox{if }X_{t}=1;\\ f_{1}(t)+U_{t}\cdot f_{0}(t)&\mbox{if }X_{t}=0,\end{cases}$ (32)
where $f_{k}(t)=t^{-1}\cdot\#\{i\leq t:X_{i}=k\}$ denotes the empirical frequency of seeing $k\in\{0,1\}$ up to time $t$, and $U_{t}\sim\mathsf{Unif}[0,1]$ is an IID Uniform random variable that is independent of $\mathbb{F}$. Each $s_{t}$ captures the “strangeness” of $X_{t}$ relative to the summary statistics (the empirical frequencies) of the data up to time $t$, along with external randomization. Equation (31) is a general definition for conformal p-values, while equation (32) represents a simplification assuming that the nonconformity score is the identity function and that the data is binary. By Vovk et al., (2003); Vovk et al., 2022a , the conformal p-values $(s_{t})_{t\geq 1}$ are IID with $\mathsf{Unif}[0,1]$ under the null hypothesis $\mathcal{H}_{0}^{\mathsf{IID}}$. Despite its name, the conformal p-value is not a p-value in the traditional sense, as it can approach either 0 or 1 under changepoint alternatives.
@@
In practice, we may “hedge our bets” across different values of $\lambda$, by taking the weighted average across (say) $\lambda\in\{-1,0,1\}$, and further “rebalance” a fraction of our portfolio at each round. These strategies yield Vovk et al., (2021)’s simple jumper in its full form:
  1. 1.
    Fix a “jump” hyperparameter $\epsilon\in(0,1)$ and normalized weights $w_{\lambda}$ for $\lambda\in\{-1,0,1\}$.
  2. 2.
    For each $\lambda\in\{-1,0,1\}$, let $\mathfrak{e}_{0}(\lambda)=w_{\lambda}$. Let $\mathfrak{e}_{0}^{\mathsf{jump}}=\mathfrak{e}_{0}(-1)+\mathfrak{e}_{0}(0)+\mathfrak{e}_{0}(1)$.
  3. 3.
    For rounds $t=1,2,\dotsc$:
    1. (a)
      Observe the new data point $X_{t}$ and compute the conformal p-value $s_{t}$.
    2. (b)
      For each $\lambda\in\{-1,0,1\}$, let $\mathfrak{e}_{t}(\lambda)=(1-\epsilon)\cdot\mathfrak{e}_{t-1}(\lambda)+\epsilon\cdot w_{\lambda}\mathfrak{e}_{t}^{\mathsf{jump}}$.
    3. (c)
      For each $\lambda\in\{-1,0,1\}$, update $\mathfrak{e}_{t}(\lambda)=\mathfrak{e}_{t}(\lambda)\cdot\left[1+\lambda\left(s_{t}-\frac{1}{2}\right)\right]$.
    4. (d)
      Let $\mathfrak{e}_{t}^{\mathsf{jump}}=\mathfrak{e}_{t}(-1)+\mathfrak{e}_{t}(0)+\mathfrak{e}_{t}(1)$.
The resulting process, $(\mathfrak{e}_{t}^{\mathsf{jump}})_{t\geq 0}$, is a test martingale for $\mathcal{H}_{0}^{\mathsf{IID}}$ in the coarse filtration $\mathbb{G}$. The conformal test martingale (33) is a special case where all weights are placed on a single choice of $\lambda$, in which case the hyperparameter $\epsilon$ becomes irrelevant. For example, the one we used in Example 1 corresponds to the simple jumper with weights $w_{1}=1$ and $w_{0}=w_{-1}=0$.
Empirically speaking, when we instead use $w_{-1}=w_{1}=1/2$ ($w_{0}=0$), with $\epsilon=0.01$, the mean of stopped e-values (as in right plot of Figure 1) also exceeds 1 when the stopping time is the first time when $k$ consecutive zeros or ones are observed, although the standard error over repeated runs is substantially larger. We find that Vovk, 2021b ’s original variant, with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$, does not empirically violate the stopping time validity in the data filtration $\mathbb{F}$, despite not having $\mathbb{F}$-anytime-validity in theory. For the simulations under the alternative hypothesis (Figures 5 and 9), and for the real data experiments (Section 4.2), we use the original version with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$.
In this paper, the specific choice of the simple jumper conformal test martingale is made solely for the clarity of exposition, and the naïve simple jumper algorithm is by no means the optimal version of the conformal test martingale. There exist other variants of conformal test martingales, including ones powerful against other types of alternatives; see Vovk et al., 2022a (, section 9). Yet, as referenced in the main text, even when combining different instantiations of conformal test martingales, the same issue of not having validity at data-dependent stopping times would still arise.
@@ B.4 Why we do not need to lift the UI e-process
Appendix C Additional example: sequentially testing a scale-invariant Gaussian mean
Here, we give another concrete example of an e-process defined in a coarser filtration $\mathbb{G}\subsetneq\mathbb{F}$ that is not valid at a $\mathbb{F}$-stopping time. This particular construction is due to Hendriksen et al., (2021); Pérez-Ortiz et al., (2024); we follow the exposition of the latter paper. The e-process described below can be combined with another e-process that is valid in $\mathbb{F}$, such as Wang and Ramdas, 2025a ’s e-process based on UI (Wasserman et al.,, 2020).
@@ Appendix C Additional example: sequentially testing a scale-invariant Gaussian mean
$\mathfrak{e}_{t}^{\mathsf{MaxInv}}=\frac{\int_{\sigma>0}\frac{1}{\sigma^{t}}\exp\left[-\frac{t}{2}\left\{\left(\frac{\hat{\mu}_{t}}{\sigma}-\delta_{1}\right)^{2}+\frac{\hat{\sigma}_{t}^{2}}{\sigma^{2}}\right\}\right]d\sigma}{\int_{\sigma>0}\frac{1}{\sigma^{t}}\exp\left[-\frac{t}{2}\left\{\left(\frac{\hat{\mu}_{t}}{\sigma}-\delta_{0}\right)^{2}+\frac{\hat{\sigma}_{t}^{2}}{\sigma^{2}}\right\}\right]d\sigma},$ (36)
@@
For this problem, we can also consider combining different e-processes in an analogous manner to the case of testing randomness (Example 1). For instance, Wang and Ramdas, 2025a (, theorem 3.2) develop a UI-based e-process for sequentially testing whether a Gaussian mean is zero under unknown variance (the “Gaussian $t$-test”), which coincides with testing the scale-invariant null $\mathcal{H}_{0}$ when $\delta_{0}=0$. (We note that, in the paper, they further discuss test martingales for the general scale-invariant null.) Analogous to Ramdas et al., (2022)’s UI e-process for testing randomness (30), and Wang and Ramdas, 2025a ’s e-process is valid in the data filtration $\mathbb{F}$. It has the following closed-form expression:
@@
To combine the two e-processes $\mathfrak{e}^{\mathsf{MaxInv}}$ (36) and $\mathfrak{e}^{\mathsf{UI}}$ (38), both of which tests whether $\mu/\sigma=0$, we can employ the adjust-then-combine strategy (Corollary 2) to obtain an e-process in $\mathbb{F}$:
@@
Appendix D The running maximum of a test martingale (often) tends to infinity in expectation, even under the null
It is known that, generally speaking, the running maximum of a test martingale or an e-process for some null hypothesis $\mathcal{P}$ is not a test martingale or an e-process for $\mathcal{P}$, respectively; see, e.g., Shafer et al., (2011, answer 1) and Ramdas et al., (2020, remark 16). To show a concrete example of when this occurs, we describe a case where we can show something even stronger: the expected running maximum under the null can tend to infinity. Then, we discuss how this represents a broad class of test martingales (and, often, e-processes).
D.1 Illustrative example: Sequentially testing a bounded IID mean
We begin with the problem of sequentially testing the mean of IID bounded random variables. Suppose that $X_{1},X_{2},\dotsc\sim P$ are IID random variables taking values in $[0,1]$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. We shall further assume that each $X_{t}$ has a finite but non-zero variance, $\sigma^{2}\in(0,\infty)$. Denoting $\mathcal{P}^{\mu}$ as all possible distributions on $[0,1]$ whose mean is $\mu$, we can test whether $P\in\mathcal{P}^{\mu}$ by constructing an e-process for the null hypothesis $\mathcal{P}^{\mu}$.
This is a composite and nonparametric testing problem that has been studied extensively in the SAVI literature. For example, Waudby-Smith and Ramdas, (2024) design powerful test martingales for $\mathcal{P}^{\mu}$ of the form
$M_{t}=\prod_{i=1}^{t}\left[1+\lambda_{i}\left(X_{i}-\mu\right)\right],\quad\forall t,$ (40)
where $(\lambda_{t})_{t\geq 1}$ is a $\mathbb{F}$-predictable sequence taking values in the interval $(0,\frac{1}{\mu}]$. (This is a special case of their “hedged capital process” where the alternative says the mean is greater than $\mu$.) The authors design various game-theoretic methods for choosing $(\lambda_{t})_{t\geq 1}$ (referred to as the betting strategy) that lead to powerful sequential tests and confidence sequences for bounded means.
For this test martingale, it is possible to derive the conditions with which the running maximum of the test martingale $(M_{t})_{t\geq 0}$ tends to infinity as $t\to\infty$, even under the null. Let $M_{t}^{*}=\sup_{i\leq t}M_{i}$ denote the running maximum of the test martingale. The result below follows from an anti-concentration lemma due to Ramdas et al., (2020).
Proposition 1.
Suppose that the $\mathbb{F}$-predictable sequence $(\lambda_{t})_{t\geq 0}$ are chosen such that, under each $P\in\mathcal{P}^{\mu}$, the sum of quadratic variations in the data, weighted by $\lambda_{t}^{2}$, tends to infinity:
$\sum_{t=1}^{\infty}\lambda_{t}^{2}\left(X_{t}-\mu\right)^{2}=\infty,\quad P\text{-a.s.}$ (41)
Then, for each $P\in\mathcal{P}^{\mu}$,
$\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]\to\infty\quad\text{as}\quad t\to\infty.$ (42)
Knowing that each $X_{t}$ has finite and non-zero variance $\sigma^{2}$, we can ensure condition (41) in different ways. The simplest case is when the predictable sequence is constant, i.e., $\lambda_{t}=\lambda\in(0,\frac{1}{\mu}]$. By the strong law of large numbers (SLLN), $\frac{1}{t}\sum_{i=1}^{t}(X_{i}-\mu)^{2}$ converges to $\sigma^{2}>0$, $P$-almost surely, so the sum $\sum_{i=1}^{t}(X_{i}-\mu)^{2}$ grows linearly in $t$ (diverging to infinity).
A practically relevant case is when the sequence decays at a rate of $\left({\sqrt{t\log t}}\right)^{-1}$ (or a slower rate). For example, Waudby-Smith and Ramdas, (2024) recommends using $\hat{\lambda}_{t}\propto\left(\sqrt{\hat{\sigma}_{t-1}^{2}t\log(t+1)}\right)^{-1}$, where $\hat{\sigma}_{t-1}^{2}$ is a $\mathcal{F}_{t-1}$-measurable, consistent estimate of the variance $\sigma^{2}$. In such cases, we may use classical generalizations of the SLLN to martingales involving predictable weights (see, e.g., theorem 2.18 of Hall and Heyde,, 1980) to show that condition (41) is met.
Proof of Proposition 1.
Consider any $P\in\mathcal{P}^{\mu}$. Denote the multiplicative increments of the test martingale as
$Y_{t}=\frac{M_{t}}{M_{t-1}}=1+\lambda_{t}\left(X_{t}-\mu\right),\quad\forall t\geq 1,$ (43)
where $0/0:=0$. Observe that $\mathbb{E}_{{P}}\left[{Y_{t}\mid\mathcal{F}_{t-1}}\right]=1$ for each $t$. We may rewrite condition (41) as
$\sum_{t=1}^{\infty}\left(Y_{t}-1\right)^{2}=\infty,\quad P\text{-a.s.}$ (44)
We also know, by construction, each $Y_{t}$ is bounded by the constant $2$. Writing $\epsilon=1$, we have
$P\left(Y_{t}\leq 1+\epsilon\right)=1.$ (45)
These two conditions, (44) and (45), are sufficient for an “anti-concentration” lemma for test martingales (Ramdas et al.,, 2020, lemma 33). In particular, the conditions imply that
$P\left(\sup_{t}M_{t}\geq x\right)\geq\frac{1}{(1+\epsilon)x}=\frac{1}{2x},\quad\forall x\in(1,\infty).$ (46)
This is enough to show that the running maximum process $M_{t}^{*}=\sup_{i\leq t}M_{i}$ tends to infinity in expectation. More precisely,
$\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]\geq\mathbb{E}_{{P}}\left[{\lim_{t\to\infty}M_{t}^{*}}\right]=\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]=\int_{1}^{\infty}P\left(\sup_{t}M_{t}\geq x\right)dx=\int_{1}^{\infty}\frac{dx}{2x}=\infty,$ (47)
where we used Fatou’s lemma and the integral formula for the expectation of non-negative random variables. (All limits exist in the extended reals.) ∎
D.2 When does this happen for test martingales (and e-processes)?
This phenomenon is not specific to the bounded IID mean testing problem—in fact, we may apply the anti-concentration lemma to any test martingale whose sum of quadratic variations is infinite. Many test martingales used in practice will satisfy this requirement: game-theoretically, as long as one does not decrease one’s bets “too quickly” (or stop betting entirely), which is the case for many powerful test martingales and e-processes, the variance in data will force the sum to infinity. Note that, for test martingales having the form (40), the bets correspond to the predictable sequence $(\lambda_{t})_{t\geq 1}$. More generally for e-processes, we remark that e-processes for composite nulls $\mathcal{P}$ are often constructed as the minimum of powerful test martingales for each null distribution $P\in\mathcal{P}$ (e.g., the UI e-process of Wasserman et al.,, 2020).
The following proof sketch is due to Martin Larsson (via personal communication). Suppose that $(M_{t})_{t\geq 0}$ is a test martingale for $P$. For $t\geq 1$, let $Y_{t}=M_{t}/M_{t-1}$ denote the multiplicative increment of this martingale ($0/0:=0$), and assume that
$\sum_{t=1}^{\infty}\left(Y_{t}-1\right)^{2}=\infty,\quad P\text{-a.s.}$ (48)
By the first half of the anti-concentration lemma (Ramdas et al.,, 2020, lemma 33), we know
$M_{\infty}=\limsup_{t\to\infty}M_{t}=0,\quad P\text{-a.s.}$ (49)
Now, suppose for contradiction that $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ were finite. Denoting $M_{\infty}^{*}=\lim_{t\to\infty}M_{t}^{*}=\sup_{t}M_{t}$, we see that $(M_{t})_{t\geq 0}$ is dominated by $M_{\infty}^{*}\in L^{1}$, and thus it is uniformly integrable (see, e.g., theorem 4.6.7 of Durrett,, 2019). This then implies that the martingale $(M_{t})_{t\geq 0}$ converges to $M_{\infty}$ in $L^{1}$. In particular, $\mathbb{E}_{{P}}\left[{M_{\infty}}\right]=\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}}\right]=\mathbb{E}_{{P}}\left[{M_{0}}\right]=1$, which contradicts equation (49). Therefore, $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ must be infinite, and by Fatou’s lemma, $\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]=\infty$.
Appendix E Other types of adjusters
E.1 “Zero” adjusters can be (even more) powerful $\mathfrak{e}$-lifters